# Architecture Decoupling Is Not All You Need For Unified Multimodal Model

### v2

@@ 4.1 Experimental Setup
Datasets. We primarily use open-source image generation and understanding datasets for training, including ShareGPT-4V [3], BLIP3-o [4], and OpenSora [52] for generation, and LLaVA-OneVision-1.0 [19], Mammoth-VL [13] for understanding, resulting in 1.5M samples for each task. Since the quality of these datasets is not as high as that of Janus-Pro’s data, we incorporate 200K internal samples to align the data quality.
@@ 4.2 Evaluation on the benchmarks
Multimodal Understanding. We evaluate our model on six widely recognized benchmarks—MME [11], MMBench (1.0-EN) [25], MMVet [49], MMMU [50], POPE [21], and MMVP [35]—which together form a compact yet thorough assessment framework encompassing perception, cognition, and multimodal reasoning, with robust capability to distinguish performance differences among leading models. The result is shown in Table 3.1, within each decoupling degree, we achieve state-of-the-art results under the same training configuration, and narrow the performance gap with models employing higher decoupling degrees or different training paradigms.
Text-to-Image Generation. We follow Janus-Pro [5] and report results on widely used image generation benchmarks Geneval [12] and DPG-Bench [16]. As shown in Table 3.1, Equipping with the AIA loss regulation, we improve the performance of Janus-Pro and Emu3 and narrow the gap with models with higher decoupling degrees.
@@ 4.3 Ablation Study
Due to the slow training speed of Emu3, we primarily conduct ablation studies on Janus-Pro. As mentioned in Sec 3.3, this represents a more challenging setting, thereby better demonstrating the effectiveness of our method.
Data Quality Analysis. Since the training data of Janus-Pro is not publicly available, we first assess the quality of our own dataset to validate our method’s effectiveness. As shown in Table 3 (i.e., w/o AIA), further fine-tuning Janus-Pro with our data yields comparable performance to the original model in both understanding and generation, indicating that our data quality itself does not introduce performance gains. Note that the Emu3 results (w/o AIA) cannot be directly compared with those reported in the original Emu3 paper, as their performance was obtained through single-task SFT. The relatively low unified performance, particularly in understanding, highlights the inherent limitations of using a VAE as the image encoder and the challenges of purely unified training.
@@
The Effectiveness of AIA loss. The result in Table 3 shows that incorporating our AIA loss leads to improved performance in both understanding and generation, with earlier integration (e.g., during the Emu3 SFT stage) yielding greater improvements. Fig 5 further illustrates the changes in cross-modal attention patterns after applying the AIA loss. With AIA regularization, the attention patterns of both Emu3 and Janus-Pro shift closer to those of single-task models, confirming that aligning attention patterns toward task-specific behaviors indeed enhances model performance, and that the AIA loss effectively improves results while reshaping attention patterns. However, it could also be observed that when adding AIA loss to purly unified model architecture Emu3, the cross-modal attention pattern is more harder to change (i.e., the generation attention pattern in Emu3 merely captures the correct directional trend across layers, but the actual values are still incorrect).
@@
Then we validate the effectiveness of huber loss and stage-level intensity. Both components are designed to relax overly strict attention constraints. As shown in the table 3, removing either leads to performance below the baseline, confirming that excessively rigid attention constraints hinder training. The Huber loss and stage-level intensity provide coarse optimization targets while allowing the model flexibility for self-adjustment, effectively addressing this issue and validating the effectiveness of AIA loss.
@@
Task-specific Attention Patterns Selection. To identify the most suitable task-specific attention patterns as training targets for unified models, we compare Deepseek-VL2 [23], the InternVL [6, 53] series, and the Qwen [1] series for understanding tasks, and FLUX [18], SimpleAR [38], Qwen-Image [41] and HunyuanImage-3.0 [2] for generation tasks. Fig 4 illustrates the corresponding attention patterns of these models. We observe that attention patterns for understanding tasks remain highly consistent across different models, whereas those for generation tasks exhibit significant variations. Consequently, we select Qwen3-VL-8B as the target for understanding and combine it with attention patterns from four generation models. Table 3 shows that the attention pattern from HunyuanImage-3.0 achieves the best performance as a learning target, while that from Qwen-Image yields mediocre results. We attribute this to differences in attention mechanisms: Qwen-Image employs a fixed MLLM module, limiting us to extracting full attention patterns from its diffusion head, which are not well-suited for autoregressive models. However, we also argue that HunyuanImage-3.0 may not represent the optimal choice, as its training incorporates understanding tasks and undergoes reinforcement learning, making its attention patterns potentially suboptimal for the PT or SFT stages of unified model training.
@@
$\lambda$ Selection. Here, we analyze the coefficient of the AIA loss during fine-tuning on Janus-Pro. We show the results in Table 3, where NTP:AIA denotes the loss weight ratio between the two objectives. By modifying attention patterns in a model with well-established pretrained knowledge is highly sensitive to the loss coefficient (i.e., also validated in Fig 3). The coefficient must be carefully balanced—strong enough to influence model optimization, yet not so dominant as to disrupt the existing knowledge.
@@
Is Data Sampling Ratios still Matter with AIA loss? Previous methods, beyond architecture decoupling, primarily mitigate task conflicts by adjusting the ratio of understanding and generation data. As concluded in BAGEL, understanding tasks converge faster than generation tasks, leading to a data distribution heavily skewed toward generation in later training stages. In this work, we alleviate task conflicts through the AIA loss and further investigate how this affects the model’s sensitivity to data sampling ratios. As shown in Table 3, we find that a balanced 1:1 ratio achieves the best performance, contrary to the conventional high-generation, low-understanding distribution. This suggests that with the AIA loss, understanding and generation tasks not only experience reduced conflict but also exhibit synergistic effects (i.e., training with both tasks outperforms using generation data alone). We believe this represents a significant step forward in unified model training.
@@ 4.4 Discussions
The Task-Specific Attention Patterns Difference in Various Models. As shown in Fig 4, for understanding tasks, models across different series and sizes exhibit nearly identical cross-modal attention patterns due to their shared autoregressive architecture, reflecting the maturity of current understanding architectures. In contrast, generation tasks present diverse patterns due to two distinct architectural paradigms: autoregressive with diffusion head and pure diffusion. Notably, models within the same architectural type display consistent attention patterns. We further observe that generation attention patterns closely follow their underlying attention mechanisms: models with bidirectional attention (FLUX, Qwen-Image) maintain consistent cross-modal interaction intensity across all layers, while those with causal attention (HunyuanImage-3.0, SimpleAR) progressively reduce attention to text as image tokens are generated—completely different from understanding tasks. Interestingly, even with the same attention mechanism, understanding and generation tasks exhibit distinct cross-modal patterns. This difference may hold the key to resolving task conflicts in unified training.
@@
What is the Right Path toward Unified Models? Through our investigation, we find that regardless of architecture decoupling strategies, generation and understanding tasks consistently exhibit mutually exclusive cross-modal interactions within the same network layers. This aligns with our earlier observation that even under the same autoregressive architecture, understanding and generation display distinct cross-modal patterns when trained independently, suggesting that the task conflicts transcend s any architecture design. However, this raises an intriguing question: if models invariably learn task-exclusive interaction patterns during unified training—regardless of architecture choices—could this actually represent the correct behavior for unified models? Despite the negative correlation between tasks, models can identify the current task through input sequences (image-text or text-image) and special tokens (e.g., <<img start>>), and automatically adjust cross-modal interaction accordingly. With appropriate explicit guidance methods like our AIA, task conflicts may not be an issue to avoid, but rather a natural characteristic to manage.
An alternative path toward unification would be to remove all task-distinguishing cues (i.e., adopt a unified tokenizer, eliminate task-related special tokens, and use interleaved data to obscure input formats), forcing the model to learn a truly unified space from the input alone. While this approach may resolve negative correlations between tasks, it would significantly increase training difficulty. We leave this exploration to future work.
@@ 5 Conclusion
Supplementary Material
In this supplementary file, we provide comprehensive details of the layer-wise target boundary $T_{l}$ and Huber threshold $\delta_{l}$ in Sec F.
F Details on layer-wise target boundary $T_{l}$ and Huber threshold $\delta_{l}$
Here we show the detailed hyper-parameters of Huber loss in Emu3 and Janus-Pro in Table S4 and Table S5 respectively. We use nearly identical parameters for each task across both models to enhance reliability. Note that the slight differences in the layer ranges of $l$ are due to the different number of layers in Emu3 and Janus-Pro.
Table S4: The hyper-parameter of Huber loss in Emu3.
$0\leq l<10$ $(0.2,0.4)$ $(0.05,0.1)$
$10\leq l<20$ $(0.1,0.4)$ $(0.05,0.15)$
$20\leq l<25$ $(0.1,0.4)$ $(0.05,0.3)$
$25\leq l\leq 30$ $(0.05,0.2)$ $(0.05,0.3)$
$l>30$ $(0.05,0.2)$ $(0.05,0.2)$
Table S5: The hyper-parameter of Huber loss in Janus-Pro.
$0\leq l<10$ $(0.2,0.4)$ $(0.05,0.1)$
$10\leq l<20$ $(0.1,0.4)$ $(0.05,0.15)$
$20\leq l<25$ $(0.1,0.4)$ $(0.05,0.3)$
$25\leq l\leq 29$ $(0.05,0.2)$ $(0.05,0.3)$
$l>29$ $(0.05,0.2)$ $(0.05,0.2)$

### v3

@@ 4.1 Experimental Setup
Datasets. We primarily use open-source image generation and understanding datasets for training, including ShareGPT-4V Chen et al. (2025b), BLIP3-o Chen et al. (2025a), and OpenSora Zheng et al. (2024) for generation, and LLaVA-OneVision-1.0 Li et al. (2024), Mammoth-VL Guo et al. (2024) for understanding, resulting in 1.5M samples for each task. Since the quality of these datasets is not as high as that of Janus-Pro’s data, we incorporate 200K internal samples to align the data quality.
@@ 4.2 Evaluation on the benchmarks
Multimodal Understanding. We evaluate our model on six widely recognized benchmarks—MME Fu et al. (2023), MMBench (1.0-EN) Liu et al. (2024b), MMVet Yu et al. (2023), MMMU Yue et al. (2024), POPE Li et al. (2023), and MMVP Tong et al. (2024)—which together form a compact yet thorough assessment framework encompassing perception, cognition, and multimodal reasoning, with robust capability to distinguish performance differences among leading models. The result is shown in Table˜3, within each decoupling degree, we achieve state-of-the-art results under the same training configuration, and narrow the performance gap with models employing higher decoupling degrees or different training paradigms.
Text-to-Image Generation. We follow Janus-Pro Chen et al. (2025c) and report results on widely used image generation benchmarks Geneval Ghosh et al. (2023) and DPG-Bench Hu et al. (2024). As shown in Table˜3, equipping with the AIA loss regulation, we improve the performance of Janus-Pro and Emu3 and narrow the gap with models with higher decoupling degrees.
Table 5: Quantitative comparison on Real-Unify.
Method Janus-Pro Janus-Pro+AIA
Step-wise results 25.2 27.8
Interleave Reasoning. Here we show the improvement on interleave reasoning ability. Since current fully unified models and encoder-decoupled methods are still in the early stages of development and have not yet demonstrated mature interleaved reasoning abilities (i.e., evidenced by the fact that no experiments regarding these models are reported in the interleaved reasoning benchmark Uni-MMMU Zou et al. (2025)). We primarily validate our method on the “Understanding aids Generation” metric (which serves as a step toward interleaved reasoning) from RealUnify Shi et al. (2025) on Janus-Pro. The result in Table˜5 shows that AIA improves the interleave reasoning performance compared with Janus-Pro.
@@ 4.3 Ablation Study
Due to the slow training speed of Emu3, we primarily conduct ablation studies on Janus-Pro. As mentioned in section˜3.3, this represents a more challenging setting, thereby better demonstrating the effectiveness of our method.
Data Quality Analysis. Since the training data of Janus-Pro is not publicly available, we first assess the quality of our own dataset to validate our method’s effectiveness. As shown in Table˜4 (i.e., w/o AIA), further fine-tuning Janus-Pro with our data yields comparable performance to the original model in both understanding and generation, indicating that our data quality itself does not introduce performance gains. Note that the Emu3 results (w/o AIA) cannot be directly compared with those reported in the original Emu3 paper, as their performance was obtained through single-task SFT. The relatively low unified performance, particularly in understanding, highlights the inherent limitations of using a VAE as the image encoder and the challenges of purely unified training.
@@
The Effectiveness of AIA loss. The result in Table˜4 shows that incorporating our AIA loss leads to improved performance in both understanding and generation, with earlier integration (e.g., during the Emu3 SFT stage) yielding greater improvements. fig.˜4 further illustrates the changes in cross-modal attention patterns after applying the AIA loss. With AIA regularization, the attention patterns of both Emu3 and Janus-Pro shift closer to those of task-specific models, confirming that aligning attention patterns toward task-specific behaviors indeed enhances model performance, and that the AIA loss effectively improves results while reshaping attention patterns. However, it could also be observed that when adding AIA loss to purly unified model architecture Emu3, the cross-modal attention pattern is more harder to change (i.e., the generation attention pattern in Emu3 merely captures the correct directional trend across layers, but the actual values are still incorrect).
@@
Then we validate the effectiveness of huber loss and stage-level intensity. Both components are designed to relax overly strict attention constraints. As demonstrated in Table˜4, eliminating either component results in performance degradation below the baseline, suggesting that overly rigid attention constraints impede the training process. The Huber loss and stage-level intensity provide coarse optimization targets while allowing the model flexibility for self-adjustment, effectively addressing this issue and validating the effectiveness of AIA loss.
Performance Gain From Knowledge Distillation or AIA? Although we rely on coarse-grained cross-modal attention statistics from task-specific models (see section˜3.1 for calculation details), a fundamental question arises regarding the source of our gains. We need to determine whether the performance improvement stems from general knowledge distillation or specifically from the alignment of attention patterns. This distinction is crucial for verifying that attention interaction patterns are indeed the primary cause of performance disparities across different architectures. To investigate this, Table˜4 compares our AIA method against standard feature-based and output-based distillation on Janus-Pro.
The results indicate that naively applying traditional knowledge distillation fails to improve performance in UMM training. This is due to the inherent conflict between generation and understanding tasks, which persists even during distillation. In contrast, constraining attention interaction patterns is the only effective strategy for performance gains. This finding validates our motivation that architecture decoupling improves performance by shifting attention patterns.
@@
Task-specific Attention Patterns Selection. To identify the most suitable task-specific attention patterns as training targets for unified models, we compare Deepseek-VL2 Liu et al. (2024a), the InternVL Chen et al. (2024); Zhu et al. (2025) series, and the Qwen Bai et al. (2025b, a) series for understanding tasks, and FLUX Labs (2024), SimpleAR Wang et al. (2025b), Qwen-Image Wu et al. (2025a) and HunyuanImage-3.0 Cao et al. (2025) for generation tasks. fig.˜5 illustrates the corresponding attention patterns of these models. We observe that attention patterns for understanding tasks remain highly consistent across different models, whereas those for generation tasks exhibit significant variations. Consequently, we select Qwen3-VL-8B as the target for understanding and combine it with attention patterns from four generation models. Table˜4 shows that the attention pattern from HunyuanImage-3.0 achieves the best performance as a learning target, while that from SimpleAR and FLUX yield mediocre results. This suggests that the quality of the generated attention patterns depends on the model’s inherent capacity, rather than the choice of attention mechanism (causal vs. bidirectional). We also argue that HunyuanImage-3.0 may not represent the optimal choice, as its training undergoes reinforcement learning, making its attention patterns potentially suboptimal for the PT or SFT stages of unified model training.
@@
$\lambda$ Selection. Here, we analyze the coefficient of the AIA loss during fine-tuning on Janus-Pro. We show the results in Table˜4, where NTP:AIA denotes the loss weight ratio between the two objectives. By modifying attention patterns in a model with well-established pretrained knowledge is highly sensitive to the loss coefficient (i.e., also validated in fig.˜3). The coefficient must be carefully balanced—strong enough to influence model optimization, yet not so dominant as to disrupt the existing knowledge.
@@
Is Data Sampling Ratios still Matter with AIA loss? Previous methods, beyond architecture decoupling, primarily mitigate task conflicts by adjusting the ratio of understanding and generation data. As concluded in BAGEL Deng et al. (2025), understanding tasks converge faster than generation tasks, leading to a data distribution heavily skewed toward generation in later training stages. In this work, we alleviate task conflicts through the AIA loss and further investigate how this affects the model’s sensitivity to data sampling ratios. As shown in Table˜4, we find that a balanced 1:1 ratio achieves the best performance, contrary to the conventional high-generation, low-understanding distribution. This suggests that with the AIA loss, understanding and generation tasks not only experience reduced conflict but also exhibit synergistic effects (i.e., training with both tasks outperforms using generation data alone). We believe this represents a significant step forward in unified model training.
Generality across Model and Data Scales. While our main experiments focus on the 7B model with 3M data, we further investigate the generality of our approach across different model sizes and data regimes. As shown in Table˜6, our method outperforms the baseline in all settings. Notably, even in low-resource scenarios, our approach yields significant gains, demonstrating its robustness and data efficiency. This confirms that AIA is not limited to large-scale training but is a versatile strategy applicable to various computational budgets.
Table 6: Generality Analysis. We compare the performance of the baseline and our method across different model sizes (1B, 7B) and training data scales (600k, 3M).
Data Size Method Image Understanding Image Generation
MMMU MMBench MMVP MMVet POPE MME-P GenEval DPG
Data-600k
1B Model Janus-Pro 36.1 61.2 43.1 37.4 86.5 1448.0 0.72 82.93
w AIA 37.2 65.6 43.8 38.5 87.1 1467.4 0.73 83.21
7B Model Janus-Pro 40.1 70.4 47.3 48.4 87.5 1587.4 0.80 84.06
w AIA 41.1 73.7 48.7 49.0 88.9 1637.8 0.80 84.78
Data-3M
1B Model Janus-Pro 36.6 62.1 42.4 38.9 86.7 1424.0 0.73 83.01
w AIA 37.7 67.4 43.8 39.4 87.6 1487 0.75 83.67
7B Model Janus-Pro 40.7 71.5 47.3 49.2 88.1 1593.1 0.80 84.19
w AIA 42.1 75.6 48.0 49.8 89.8 1656.4 0.81 84.49
@@ 4.4 Discussions
The Task-Specific Attention Patterns Difference in Various Models. As shown in fig.˜5, for understanding tasks, models across different series and sizes exhibit nearly identical cross-modal attention patterns due to their shared autoregressive architecture, reflecting the maturity of current understanding architectures. In contrast, generation tasks present diverse patterns due to two distinct architectural paradigms: autoregressive with diffusion head and pure diffusion. Notably, models within the same architectural type display consistent attention patterns. We further observe that generation attention patterns closely follow their underlying attention mechanisms: models with bidirectional attention (FLUX, Qwen-Image, HunyuanImage-3.0) maintain consistent cross-modal interaction intensity across all layers, while those with causal attention (SimpleAR) progressively reduce attention to text as image tokens are generated—completely different from understanding tasks. Interestingly, even with the same attention mechanism, understanding and generation tasks exhibit distinct cross-modal patterns. This difference may hold the key to resolving task conflicts in unified training.
@@
What is the Right Path toward Unified Models? Through our investigation, we find that regardless of architecture decoupling strategies, generation and understanding tasks consistently exhibit mutually exclusive cross-modal interactions within the same network layers. This aligns with our earlier observation that even under the same autoregressive architecture, understanding and generation display distinct cross-modal patterns when trained independently, suggesting that the task conflicts transcends any architecture design. However, this raises an intriguing question: if models invariably learn task-exclusive interaction patterns during unified training—regardless of architecture choices—could this actually represent the correct behavior for unified models? Despite the negative correlation between tasks, models can identify the current task through special tokens (e.g., <<img start>>), and automatically adjust cross-modal interaction accordingly. With appropriate explicit guidance methods like our AIA, task conflicts may not be an issue to avoid, but rather a natural characteristic to manage.
An alternative path toward unification would be unified models with a diffusion head. The core conflict between generation and understanding stems from their divergent output requirements. Understanding tasks demand high-level semantic information, whereas generation necessitates the reconstruction of fine-grained details Yu et al. (2025). Even with a unified tokenizer Tang et al. (2025); Qu et al. (2025); Yao et al. (2025); Ma et al. (2025); Lu et al. (2025), this fundamental tension persists. To resolve this, we propose aligning the information density of image and text outputs within the unified model (i.e., this is also mentioned in RecA Xie et al. (2025a)). Specifically, the model generates only abstract conceptual images, which then serve as conditions for the diffusion head to restore fine-grained details. This approach effectively eliminates the conflict between generation and understanding within the unified framework, which is a more realistic way toward real UMM and we take this as future works.